OpenStudy (kt7779):

need to turn into devision i give medals! 10=x%*40

OpenStudy (anonymous):

given, 10=x%*40 =>10=(x/100)*40 =>10=40x/100 =>1000=40x =>1000/40=x =>x=100/4=25